# GroupNL: Low-Resource and Robust CNN Design over Cloud and Device

### v1

@@ 3.3.3 Selection of NLFs, Hyperparameters and Coefficients
Nonlinear Transformation Formulation Hyperparameters
@@
In this case, we can observe that the GroupNL model with Sinusoidal function outperforms the other three NLFs and also has a similar performance to the vanilla CNN model. As a result, we can generate feature maps by the Sinusoidal function as the NLF in subsequent experiments.
Figure 4: Comparison of different NLFs in GroupNL ResNet-18 on the CIFAR-10 dataset.
Analysis of Sparse Seed Filters in GroupNL: As described in Section 3.3.1, we designate that the seed filters ${\bm{W}}_{\mathrm{seed}}$ in basic structure (e.g., ResNet-18) utilize the standard convolution, and the seed filters in bottleneck structure (e.g., ResNet-101) are sparse with convolutional groups of $\mathrm{Gcd}(c_{\mathrm{in}},c_{\mathrm{seed}})$. The sparse seed filter experimental results are shown in Table 3. Applying sparsity to the seeded convolution of ResNet-18 can introduce an extremely low number of parameters of 0.05M and just approximately 7M FLOPs. However, performance is severely degraded. As a comparison, the ResNet-101 with sparsity can maintain the performances with the lowest number of parameters and FLOPs. Therefore, the sparse seed filters are not used in the basic structure.
TABLE 3: The ablation studies of w/ and w/o sparse seed filters and different reduction ratio $r$ in GroupNL.
Model $g$ $r$Params (M) #FLOPs (M) Top-1 Acc (%)
@@
Hyperparameters and Coefficients of GroupNL: We designate that the hyperparameters of NLFs are randomly initialized and non-learnable during model training. This trick reduces the number of gradients for optimization in training while preventing the NLFs lose the data-agnostic properties. The Sinusoidal function has two hyperparameters $\omega$ and $\phi$, we randomly choose them via $\tt{torch.rand}$ between $[1,2]$ and $[1,5]$, respectively.
@@
Besides, for the ResNet-101, the number of parameters and FLOPs are very similar for different $r$ settings. This is due to the bottleneck structure in ResNet-101 using the sparse seed filters with convolutional groups of $\mathrm{Gcd}(c_{\mathrm{in}},c_{\mathrm{seed}})$. For example, assuming a $k^{2}$ convolution with same input and output channels $c_{\mathrm{in}}=c_{\mathrm{out}}=512$ and feature maps size $wh$ in the bottleneck. In the case of $r=2$, the number of parameters is $\mathcal{P}_{r=2}=\frac{c_{\mathrm{in}}c_{\mathrm{seed}}k^{2}}{\mathrm{Gcd}(c_{\mathrm{in}},c_{\mathrm{seed}})}=\frac{512\cdot 256k^{2}}{256}=512k^{2}$, and the FLOPs is $\mathcal{F}_{r=2}=\frac{c_{\mathrm{in}}c_{\mathrm{seed}}k^{2}wh}{\mathrm{Gcd}(c_{\mathrm{in}},c_{\mathrm{seed}})}=\frac{512\cdot 256k^{2}wh}{256}=512k^{2}wh$. In the cases of $r=\{4,16,32,\cdots,256,512\}$, the results are also $\mathcal{P}_{r}=512k^{2}$ and $\mathcal{F}_{r}=512k^{2}wh$ because of the fraction $c_{\mathrm{seed}}$ are reduced in the case of using the greatest common divisor grouping seed filters.
@@
TABLE 4: The ablation studies of different splitting groups $g$ in GroupNL CNN models.
Model $r$ $g$Params (M) #FLOPs (M) Top-1 Acc (%)
@@
3.4 Resource and Training Acceleration Analysis
For cloud-assisted CNN models, the multi-GPUs training on cloud servers is mainly based on data parallelism. In practical multi-GPU training, the bandwidth of inter-GPU memory data transfer is an important constraint on the training speed compared to the powerful and fast computation kernels. In the classic data parallelism, e.g., $\tt{torch.nn.DataParallel}$, the data exchange between different GPUs in forward/backward of the model generally in three stages: (i) Gather outputs (or logits) and compute loss gradients on host GPU; (ii) Scatter gradients to multi-GPUs; (iii) Reduce gradients to host GPUs. For the distributed data parallelism, e.g., $\tt{torch.nn.parallel.DistributedDataParallel}$, only one reduction of gradients between GPUs in each backward.
So, the number of gradients can be further optimized in terms of inter-GPU scatters and reductions in different data parallelism. Assuming each trainable parameter in a CNN model maintains a gradient, the number of gradients $\mathcal{G}$ corresponds to the number of trainable parameters $\mathcal{P}$.
At the same time, designing libraries-agnostic and task-specific CNN methods is the minimal change in cloud-assisted architecture for generalization. Based on this premise, we analyze the theoretical gradient exchange in multi-GPUs memory of vanilla CNN, GhostNet, SineFM, and GroupNL in training.
Ignoring the bias, for a vanilla convolutional layer, the number of parameters and gradients can be described as follows:
$\begin{array}[]{l}{\mathcal{P}_{\mathrm{vanilla\ cnn}}=\mathcal{G}}_{\mathrm{vanilla\ cnn}}=c_{\mathrm{in}}c_{\mathrm{out}}k^{2}\ .\end{array}$ (13)
For a GhostNet layer, the number of parameters and gradients from the seed filters and $d\times d$ DWConv are:
${\mathcal{P}_{\mathrm{ghostnet}}=\mathcal{G}}_{\mathrm{ghostnet}}$ $=c_{\mathrm{in}}c_{\mathrm{seed}}k^{2}+c_{\mathrm{gen}}d^{2}$ (14)
$=c_{\mathrm{in}}c_{\mathrm{out}}k^{2}/r+c_{\mathrm{out}}(1\!-\!\frac{1}{r})d^{2}\ ,$
and the number of parameters and gradients for the SineFM layer from the seed filters, $1\times 1$ DWConv, and batch normalization are:
${\mathcal{P}_{\mathrm{sinefm}}=\mathcal{G}}_{\mathrm{sinefm}}$ $=c_{\mathrm{in}}c_{\mathrm{seed}}k^{2}+tc_{\mathrm{gen}}\cdot 1^{2}+2tc_{\mathrm{seed}}$ (15)
$=c_{\mathrm{in}}c_{\mathrm{out}}k^{2}/r+tc_{\mathrm{out}}(1\!-\!\frac{1}{r})+2tc_{\mathrm{out}}/r$
$=c_{\mathrm{in}}c_{\mathrm{out}}k^{2}/r+c_{\mathrm{out}}(1\!+\!\frac{1}{r})t\ .$
The proposed GroupNL layer only contains the learnable standard seed filters and learnable sparse seed filters, and the number of parameters and gradients are shown in Eq. 16 and Eq. 17, respectively:

### v2

@@ 4.4.1 Selection of Nonlinear Functions
Nonlinear Formulation Hyperparameters
@@
In this case, we can observe that the GroupNL model with the Sinusoidal function outperforms the other three NLFs and also has a similar performance to the vanilla CNN model. As a result, in our experiments, we use the Sinusoidal function as the NLF to generate feature maps.
We offer a conjecture: due to the periodicity of the Sinusoidal function and its different frequencies and phase hyperparameters, it can generate a richer variety of “similar but phase-different patterns” compared to other functions, which meets the requirement of GroupNL to generate similar but diverse feature maps. The reasons why other functions are worse may be: Monomial only performs power transformations and does not generate new spatial patterns, resulting in insufficient diversity; Gaussian function has too strong locality, with values only near the function center, and other values are suppressed to 0, which cannot guarantee the similarity of the generated images; Laplace function is similar to Gaussian.
Figure 11: Comparison of different NLFs in GroupNL and vanilla ResNet-18 on the CIFAR-10 dataset.
4.4.2 Selection of Hyperparameters and Coefficients
Analysis of Sparse Seed Filters in GroupNL: As described in Section 3.2.1, we designate that the seed filters ${\bm{W}}_{\mathrm{seed}}$ in basic structure (e.g., ResNet-18) utilize the standard convolution, and the seed filters in bottleneck structure (e.g., ResNet-101) are sparse with convolutional groups of $\mathrm{Gcd}(c_{\mathrm{in}},c_{\mathrm{seed}})$. The sparse seed filter experimental results are shown in Table 14. Applying sparsity to the seeded convolution of ResNet-18 can introduce an extremely low number of parameters of 0.05M and approximately 7M FLOPs. However, performance is severely degraded. As a comparison, the ResNet-101 with sparsity can maintain the performance with the lowest number of parameters and FLOPs. Therefore, the sparse seed filters are not used in the basic structure.
Additionally, Sparse GroupNL modules are designated to replace the depthwise convolutions in lightweight CNNs [34, 17, 8, 36, 7, 37] to maintain the compactness of these models in terms of number of parameters and FLOPs.
TABLE 14: Ablation Studies of w/ and w/o Sparse Seed Filters and Reduction Ratio $r$ in GroupNL.
Model $g$ $r$ #Params (M) #FLOPs (M) Top-1 Acc (%)
@@ 4.4.2 Selection of Hyperparameters and Coefficients
Hyperparameters and Coefficients of GroupNL: We designate that the hyperparameters of NLFs are randomly initialized and non-learnable during model training. This trick reduces the number of gradients for optimization in training while preventing the NLFs from losing the data-agnostic properties. The Sinusoidal function has two hyperparameters $\omega$ and $\phi$, we randomly choose them via $\tt{torch.rand}$ between $[1,2]$ and $[1,5]$, respectively.
@@
Besides, for the ResNet-101, the number of parameters and FLOPs are very similar for different $r$ settings. This is due to the bottleneck structure in ResNet-101 using the sparse seed filters with convolutional groups of $\xi=\mathrm{Gcd}(c_{\mathrm{in}},c_{\mathrm{seed}})$. For example, assuming a $k^{2}$ convolution with same input and output channels $c_{\mathrm{in}}\!=\!c_{\mathrm{out}}\!=\!512$ and feature maps size $wh$ in the bottleneck.
In the case of r=2r\!=\!2, the number of parameters is $\mathcal{P}_{r2}\!=\!c_{\mathrm{in}}c_{\mathrm{seed}}k^{2}\frac{1}{\xi}\!=\!512\cdot 256k^{2}\frac{1}{256}\!=\!512k^{2}$, and the FLOPs is $\mathcal{F}_{r2}\!=\!whc_{\mathrm{in}}c_{\mathrm{seed}}k^{2}\frac{1}{\xi}\!=\!wh\cdot 512\cdot 256k^{2}\frac{1}{256}\!=\!512k^{2}wh$. In the cases of $r\!=\!\{4,16,32,\cdots,256,512\}$, the results are also $\mathcal{P}_{r}\!=\!512k^{2}$ and $\mathcal{F}_{r}\!=\!512k^{2}wh$ because of the fraction $c_{\mathrm{seed}}$ are reduced in the case of using the greatest common divisor grouping seed filters.
@@
TABLE 15: Ablation Studies of Groups $g$ in GroupNL.
$r$ $g$ #Params (M) #FLOPs (M) Top-1 Acc (%)
@@
4.4.3 Module-Level Profiling
We conduct a module-level profiling of GroupNL Conv on Raspberry Pi, based on PyTorch-CPU backend. For Conv modules, the number of input and output channels is 512, the input feature maps size is (512, 32, 32), kernel size is (3, 3), stride is 2, and padding is 1, so the output feature maps size is (512, 16, 16). The FPS reports the average Frames Per Second in 100-times inference with randomly initialized inputs.